OpenStudy (anonymous):

can you halp me find the center of this circle please with this equation : x squared plus y squared plus 16x minus 16y plus 79=0

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

x^2+y^2-16x-16y+79

OpenStudy (anonymous):

complete the square

OpenStudy (anonymous):

(x^2-16x+64)+(y^2-16y+64)=-79+128 (x-8)^2+(y-8)^2=...

OpenStudy (anonymous):

adding on to what imranmeah91 already has, the center of the circle is located at the coordinates (-8,8). Although the computation that imranmeah91 has may be a little off.

OpenStudy (anonymous):

- correction x^2+y^2+16x-16y+79

OpenStudy (anonymous):

(x^2+16x+64)

OpenStudy (anonymous):

The actual equation for this particular circle is (x+8)^2+(x-8)^2=49

OpenStudy (anonymous):

again, the circle has a center at (-8,8) with a radius of 7.

OpenStudy (anonymous):

thank you